OpenStudy (anonymous):

Solve this equation: –9h – 6 + 12h + 40 = 22

OpenStudy (abmon98):

3h+34=22 3h=22-34 3h=-12 h=-4

OpenStudy (haseeb96):

-9h +12h -6 +40 -22=0 3h +12=0 3h =-12 h=-4 this is the correct answer